OpenStudy (anonymous):

pre calc simplify sec^2y-cot^2((pi/2) -y)

OpenStudy (anonymous):

\[\sec^2y-\cot^2(\frac{ \pi }{ 2 }-y)\]

OpenStudy (anonymous):

Do you use the confunction identity \[\cot (\frac{ \pi }{ 2 }-\theta)\]

OpenStudy (anonymous):

yes, the answer is 1